OpenStudy (anonymous):

the average of three cosective integer such that twice the greatest integer is 2 less than times the least integer

OpenStudy (anonymous):

@I_Need_Help_With_Home

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

Let the three consecutive numbers be : \((x-1)\), \(x\), \((x+1)\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Average of three numbers is given by: \[Average = \frac{\tt {Sum \ of \ three \ numbers}}{\tt{Number \ of \ Numbers}}\]

OpenStudy (anonymous):

right

OpenStudy (anonymous):

As you have three numbers, so their sum would be: \(Sum = x-1 + x + x+1\) implies \(Sum = 3x\)

OpenStudy (anonymous):

And as you have three numbers, so number of numbers are \(3\).

OpenStudy (anonymous):

And you have also one relation given in question, we try to use that. out of that three numbers, \((x+1)\) would be greatest of all, and \((x-1)\) will be smallest number.

OpenStudy (anonymous):

And the relation is: \(2(x+1) = 2(x-1) + 2\)

OpenStudy (anonymous):

Sorry: \(2(x+1) = 2(x-1) - 2\)

OpenStudy (anonymous):

Can you find out \(x\) from here?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Then find out \(x\), and tell me.

OpenStudy (anonymous):

less than 3 times the least integer ??

OpenStudy (anonymous):

@waterineyes the average of three cosective integer such that twice the greatest integer is 2 less than3 times the least integer

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

Then it would be: \(2(x+1) = 3(x-1)-2\)

OpenStudy (anonymous):

Can you find \(x\) here now?

OpenStudy (anonymous):

@rosebird doing?

OpenStudy (anonymous):

tag me when you are done. :)

OpenStudy (anonymous):

yeh

OpenStudy (anonymous):

why 3 befor bracket??

OpenStudy (anonymous):

why u put 3 before bracket? @waterineyes

OpenStudy (anonymous):

7? is answer :) ty @waterineyes

OpenStudy (anonymous):

3 times the least integer means 3 multiplied by least integer and least integer is \((x-1)\). So, \(3 \times (x-1)\)

OpenStudy (anonymous):

\(x =7\), but it is \(x\) value. Now we have to find Average.

OpenStudy (anonymous):

but the answers are 6 7 8

OpenStudy (anonymous):

I already calculated above that sum is \(3x\) and number of numbers you have is \(3\): So: \[Average = \frac{Sum}{Total number} = \frac{3x}{3} \implies \color{green}{Average = x}\]

OpenStudy (anonymous):

It is your luck that value of \(x\) is average itself. :P

OpenStudy (anonymous):

hahaha thanks:) @waterineyes

OpenStudy (anonymous):

Do you want to know how \(x\) is average itself?